OpenStudy (anonymous):

Solve the equation. If necessary, round your answer to the nearest tenth. x^2 = 98

OpenStudy (anonymous):

49, −49 10, −10 about 9.9, −9.9 about 6.9, −6.9

OpenStudy (anonymous):

x^2=98 \[x= \sqrt{98}\] x= 9.89 , -9.89 x=9.9 , -9.9

OpenStudy (anonymous):

@CO_oLBoY can you help me with one moree

OpenStudy (anonymous):

yeah go on

OpenStudy (anonymous):

but hurry up

OpenStudy (anonymous):

Solve the equation. 4x^2 − 8 = 56

OpenStudy (anonymous):

2, −2 4, −4 8, −8 16, −16

OpenStudy (anonymous):

4x^2 -8 = 56 4x^2 = 56 + 8 4x^2 = 64 \[x ^{2} = \frac{ 64 }{ 4 }\] x^2 = 16 x= 4 , -4

OpenStudy (anonymous):

here u go